OpenStudy (anonymous):

An elastic spring constant of 150 N / m was compressed by a block 15 cm in 5.0 kg cm. Determine the speed of the block when leaving the spring, knowing he is in a horizontal plane.

OpenStudy (anonymous):

Spring F = -kx PE = (1/2) k x^2 = (1/2)(150 N/m)(.15m)^2 = 1.6875 J PE becomes KE of block = (1/2) m v^2. Knowing PE = KE and m, find v.

OpenStudy (anonymous):

I marked the letter a) Correct?

OpenStudy (anonymous):

@douglaswinslowcooper

OpenStudy (anonymous):

1.68 J = (1/2)(5 kg)(v^2) v^2 = 0.672 v = 0.82 m/s correct. Well done.

OpenStudy (anonymous):

thank you!!!